# SemEval-2025 Task 11: Bridging the Gap in Text-Based Emotion Detection

### v1

@@ Abstract
We present our shared task on text-based emotion detection, covering more than 30 languages from seven distinct language families. These languages are predominantly low-resource and spoken across various continents. The data instances are multi-labeled into six emotional classes, with additional datasets in 11 languages annotated for emotion intensity. Participants were asked to predict labels in three tracks: (a) emotion labels in monolingual settings, (b) emotion intensity scores, and (c) emotion labels in cross-lingual settings. The task attracted over 700 participants. We received final submissions from more than 200 teams and 93 system description papers. We report baseline results, as well as findings on the best-performing systems, the most common approaches, and the most effective methods across various tracks and languages. The datasets for this task are publicly available.  [^1] The dataset is available at SemEval2024-task 11.
@@ 1 Introduction
People use language in diverse and sophisticated ways to express emotions across languages and cultures Wiebe et al. (2005); Mohammad and Kiritchenko (2018); Mohammad et al. (2018a). Emotions are also perceived subjectively, even within the same culture or social group. Recognising these emotions is central to language technologies and NLP applications in healthcare, digital humanities, dialogue systems, and beyond Mohammad et al. (2018b); Saffar et al. (2023). In this work, we use emotion recognition to refer to perceived emotions—i.e., the emotion most people believe the speaker might have felt based on a sentence or short text snippet.
@@
The task consists of three tracks: (a) multi-label emotion detection, (b) emotion intensity detection, and (c) cross-lingual emotion detection. The languages for each track are listed in Figure 1. Each team could submit results for one, two, or all three tracks in one or more languages. Our official evaluation metrics were the average F-score for Tracks A and C and the Pearson correlation coefficient for Track B, which measures how well system-predicted intensity scores align with human judgments.
Our task attracted over 700 participants, with 220 final submissions and 93 teams submitting system description papers. Track A (multi-label emotion detection) received the most submissions (114), followed by Track C (cross-lingual emotion detection) with 51, and Track B (emotion intensity detection) with 32. Most teams participated in multiple languages, averaging 11 languages per team. Our task was the most popular competition on CodaBench in 2024.  [^3] Codabench Newsletter 2024 All task details and resources are available on the task’s GitHub page. [^4] https://github.com/emotion-analysis-project/SemEval2025-Task11
@@ 2 Related Work
NLP work on emotion detection is predominantly Western-centric, with a few exceptions for languages other than English (e.g., Italian (Bianchi et al., 2021), Romanian (Ciobotaru et al., 2022), Indonesian (Saputri et al., 2018), and Bengali (Iqbal et al., 2022)). While multilingual datasets (e.g., (Öhman et al., 2020) and XLM-EMO (Bianchi et al., 2022)) exist, they do not fully capture cultural nuances in emotional expressions due to their reliance on translated data (e.g., XLM-EMO), as emotions are highly contextualized and culture-specific (Havaldar et al., 2023; Mohamed et al., 2024; Hershcovich et al., 2022). Furthermore, most datasets are single-labeled, and to the best of our knowledge, there are no multilingual resources that capture overlapping emotions and their intensity across various languages.
Additionally, most prior emotion recognition shared tasks have focused on high-resource languages such as English, Spanish, German, and Arabic Strapparava and Mihalcea (2007); Mohammad et al. (2018b); Chatterjee et al. (2019). In contrast, this shared task covers more than 30 languages, including many low-resource languages.
@@ 3.1 Data Collection
  • Literary texts: We translated the novel *La Grande Maison* (The Big House) by Mohammed Dib  [^5] https://en.wikipedia.org/wiki/La_Grande_Maison from French to Algerian Arabic (arq) and manually post-processed the translation to generate sentences to be annotated by native speakers.
  • News data: Although we prefer emotionally rich social media data from different platforms, when such data is scarce, we annotate news data and headlines in some African languages (e.g., yor, hau, and vmw).
  • Human-written and machine-generated data: We created a dataset from scratch for Hindi (hin) and Marathi (mar). We ask annotators to generate emotive sentences on a given topic (e.g., family), automatically translate a small section of the Hindi dataset to Marathi, and have native speakers manually fix translation errors. Finally, we augment both datasets with a few hundred quality-approved instances generated by ChatGPT.
@@ 3.2 Data Annotation
Once the labels for perceived emotions are assigned, the final intensity scores for Track B were determined by averaging the selected intensity values and rounding up to the nearest whole number. We only assign intensity scores for datasets where most instances were annotated by at least five annotators to ensure robustness. Table 1 shows the total number of instances in each dataset and the number of instances in the training, development, and test splits for all the languages.
@@ 4.1 Tracks
Given a text and the six emotion classes (i.e., joy, sadness, fear, anger, surprise, and disgust), participants had to predict whether the intensity of each possible emotion was equal to 0 (no emotion), 1 (low), 2 (medium), or 3 (high).
@@
Similar to Track A, participants had to predict the presence or absence of each perceived emotion, but without using any training data in the target language. Instead, participants were allowed to use labeled dataset(s) from at least one other language. For instance, one could use German data for training when testing on English.
@@ 4.2 Task Organisation
We used Codabench as the competition platform and released pilot datasets before the start of the shared task to help participants better understand the task (i.e., the datasets, the languages involved, and the labels). We provided participants with a starter kit on GitHub, resources for beginners, and organised a Q&A session along with a writing tutorial. Our participants were based in different parts of the world, as shown in Figure 2, with many coming from underrepresented regions. The task consisted of two phases: (1) the development phase and (2) the evaluation phase. During the development phase, the leaderboard was open, allowing a maximum of 999 submissions per participant. In the evaluation phase, the leaderboard was closed, and each participant was allowed up to three submissions, with the last submission being considered for the official ranking.
4.3 Evaluation Metrics and Baselines
Evaluation Metrics
Our Baselines
5 Participating Systems and Results
@@
$M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.232 $M_{\text{baseline}}$ 0.357 $M_{\text{baseline}}$ 0.2425
@@
Table 2: Average macro-F1 scores for our baselines ($M_{baseline}$ and $R_{baseline}$ referring to Majority Vote and RoBERTa baselines, respectively) and the top two performing systems for each language in Track A (in bold).
@@ 5.2.1 Best-Performing Systems
proposes one of the most effective models in the competition. They consistently rank as the top approach in Track A for 20 out of 28 languages. For their system, they combine several base models (ChatGPT-4o, DeepSeek-V3, Gemma-9b, Qwen-2.5-32b, Mistral-Small-24B) using multiple ensemble techniques (neural networks, XGBoost, LightGBM, linear regression, weighted voting). They fine-tune the Gemma and Qwen models using AdaLoRA. For prompting the LLMs, they used an iterative prompt-optimisation technique that generates prompt variations.
@@
ranks in the top 10 in 16 languages in Track A and 12th in English. They use sample contrastive learning, where performance is enhanced by comparing sample pairs, and generative contrastive learning, where the models learn to distinguish correct from incorrect predictions. Their samples are randomly selected from the task dataset (no external augmentation). They use LLaMa3-Instruct-8B for their fine-tuning.
@@ 5.2.2 Takeaways
Most of the teams that rank well on Track A experiment with essentially two methodologies: 1) fine-tuning BERT-based models such as DeBERTa, mBERT, and XLM-R; and/or 2) instruction-fine-tuning using LoRa methodologies in combination with prompt design and data augmentation techniques on LLMs (ChatGPT-4o, DeepSeek-V3, Gemma-9b, Qwen-2.5-32b, Mistral-Small-24B). For instance, Team Telai (3rd in Chinese), Team Empaths (2nd in Hausa and Somali, and 1st in Swahili), Team NYCU-NLP (2nd in English), Team JNLP (1st in Hindi and among the best four across 10 languages), Team Unibouc (1st in Emakhuwa), Team Heiderdinger (2nd in Mozambican Portuguese, German, Spanish, Brazilian Portuguese, and Yoruba; 1st in Russian), and Team Maomao (2nd in Afrikaans).
Few teams focus on only a subset of languages or explore language-related knowledge in their methodology. For instance, Team Lazarus NLP redefined and reformulated the multilabel classification into multiple binary tasks to expand training samples. They also explored how knowledge could potentially be transferred between Indonesian languages. All the top 10 teams performed much better than our baseline model, with an improvement that is more notable in a few low-resource languages, such as Oromo, where Team Tewodros obtained an average macro-F1 score of 0.616 compared to a baseline of 0.126. The same was observed for Yoruba, where Team Pai scored 0.461 compared to a baseline score of 0.092.
@@ 5.3 Track B: Emotion Intensity Detection
$R_{\text{baseline}}$ 0.405 0.562
@@
Table 3: Pearson correlation scores for our baselines (Majority $M_{baseline}$ and RoBERTa $R_{baseline}$) and the top 2 performing systems for each language in Track B (in bold).
5.3.1 Best-Performing Systems
Team Pai
Team CSECU-Learners
5.3.2 Takeaways
Teams Deepwave, Teleai, and JNLP also ranked highly across various languages using similar prompt engineering approaches as in Track A. Additionally, Team NYCU-NLP ranked second in English by aggregating instruction-tuned small language models. The different teams outperformed our RoBERTa baseline, which achieved medium Pearson correlation coefficient scores overall, with some extremely low scores in Algerian Arabic, Hausa, Ukrainian, and even Brazilian Portuguese. Overall, we observe that the teams adopted approaches similar to Track A’s, with minor adjustments in the prompts. Furthermore, we observe that even the best-performing teams achieved a Pearson correlation coefficient score of no more than 0.65 on Algerian Arabic, likely due to the novelty of the dataset.
5.4 Track C: Cross-lingual Emotion Detection
@@
$M_{\text{baseline}}$ 0.312 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.236 $M_{\text{baseline}}$
@@
$M_{\text{baseline}}$ $M_{\text{baseline}}$ 0.218 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.232
@@
$M_{\text{baseline}}$ 0.357 $M_{\text{baseline}}$ 0.243 $M_{\text{baseline}}$ 0.163 $M_{\text{baseline}}$
@@
$M_{\text{baseline}}$ 0.162 $M_{\text{baseline}}$ $M_{\text{baseline}}$ 0.165 $M_{\text{baseline}}$
Table 4: Average macro-F1 scores for our baselines ($M_{baseline}$ and $R_{baseline}$ referring to Majority Vote and RoBERTa baselines, respectively) and the top 2 performing systems for each language in Track C (in bold).
@@ 5.4.1 Best-Performing Systems
Team Deepwave fine-tuned Google Gemma-2 using tailored data augmentation and Chain-of-Thought (CoT) prompting. They decomposed the task into two sub-tasks: (1) sentiment keyword identification and (2) sentiment polarity recognition.
@@
Team maomao experimented with different setups for fine-tuning LLMs. The base models used were Qwen2.5-7B-Instruct, GPT-0544o Mini2, and LLaMA-3.2-3B-Instruct (Touvron et al., 2023). They applied Direct Preference Optimization to refine their model—a technique that selects high-quality instances from lower-quality ones within a dataset. After this step, they retrained the model using the refined dataset. They also explored random sampling and retrieval-augmented generation (RAG) methods for training, primarily on DeepSeek-V3, Qwen-Max5093, and Grok-V2.
Figure 3: Top LLMs used by participants across tracks (A, B, and C).
@@ 5.4.2 Takeaways
Other top-performing systems include Team Ozemi, who fine-tuned a multilingual BERT model and applied machine translation to enhance performance across 35 languages. They used the Synthetic Minority Oversampling Technique (SMOTE) with TF-IDF to address class imbalance in Russian. It appears they translated all datasets into a common language using Google Translate before processing, except for Nigerian Pidgin and Emakhua, where they used a multilingual BERT model for translation. They also leveraged two Kaggle competition datasets for data augmentation. Team heimerdinger, who ranked highest in Javanese, built their approach using various LLMs (LLaMA 3.1 8B, Qwen 2.5 7B, DeepSeek-7B, MistralV0.3-7B, and Gemma2-9B) for Track C. They employed in-context learning with multilingual examples from high-resource languages such as English and Spanish.
Overall, the different teams outperformed our baselines. However, we observe that the average scores for this track are lower, particularly in low-resource languages such as Emakhuwa, as participants were presented with an additional challenge.

### v2

@@ Abstract
We present our shared task on text-based emotion detection, covering more than 30 languages from seven distinct language families. These languages are predominantly low-resource and are spoken across various continents. The data instances are multi-labeled with six emotional classes, with additional datasets in 11 languages annotated for emotion intensity. Participants were asked to predict labels in three tracks: (a) multilabel emotion detection, (b) emotion intensity score detection, and (c) cross-lingual emotion detection.
The task attracted over 700 participants. We received final submissions from more than 200 teams and 93 system description papers. We report baseline results, along with findings on the best-performing systems, the most common approaches, and the most effective methods across different tracks and languages. The datasets for this task are publicly available. [^1] The dataset is available at SemEval2025 Task 11.
@@ 1 Introduction
People use language in diverse and sophisticated ways to express emotions across languages and cultures Wiebe et al. (2005); Mohammad and Kiritchenko (2018); Mohammad et al. (2018a). Emotions are also perceived subjectively, even within the same culture or social group. Recognising these emotions is central to language technologies and NLP applications in healthcare, digital humanities, dialogue systems, and beyond Mohammad et al. (2018b); Saffar et al. (2023). In this work, we use emotion recognition to refer to perceived emotions, i.e., the emotion most people believe the speaker might have felt based on a sentence or short text snippet.
@@
The task consists of three tracks: (a) multilabel emotion detection, (b) emotion intensity detection, and (c) cross-lingual emotion detection. The languages for each track are listed in Figure 1. Each team could submit results for one, two, or all three tracks in one or more languages. Our official evaluation metrics were the average F-score for Tracks A and C and the Pearson correlation coefficient for Track B, which measures how well system-predicted intensity scores align with human judgments.
Our task attracted over 700 participants, with 220 final submissions and 93 teams submitting system description papers. Track A (multi-label emotion detection) received the most submissions (114), followed by Track C (cross-lingual emotion detection) with 51, and Track B (emotion intensity detection) with 32. Most teams participated in multiple languages, averaging 11 languages per team. Our task was the most popular competition on CodaBench in 2024.333Codabench Newsletter 2024 All task details and resources are available on the task’s GitHub page. [^4] https://github.com/emotion-analysis-project/SemEval2025-Task11
@@ 2 Related Work
NLP work on emotion detection is predominantly Western-centric, with a few exceptions for languages other than English (e.g., Italian (Bianchi et al., 2021), Romanian (Ciobotaru et al., 2022), Indonesian (Saputri et al., 2018), and Bengali (Iqbal et al., 2022)). While multilingual datasets (e.g., (Öhman et al., 2020) and XLM-EMO (Bianchi et al., 2022)) exist, they do not fully capture cultural nuances in emotional expressions due to their reliance on translated data (e.g., XLM-EMO), as emotions are highly contextualized and culture-specific (Havaldar et al., 2023; Mohamed et al., 2024; Hershcovich et al., 2022). Furthermore, most datasets are single-labeled, and to the best of our knowledge, there are no multilingual resources that capture simultaneous emotions and their intensity across various languages.
Additionally, most prior emotion recognition shared tasks have focused on high-resource languages such as English, Spanish, German, and Arabic Strapparava and Mihalcea (2007); Mohammad and Bravo-Marquez (2017); Mohammad et al. (2018b); Chatterjee et al. (2019). In contrast, this shared task covers more than 30 languages, including several low-resource languages.
@@ 3.1 Data Collection
  • Literary texts: The language lead manually translated the novel La Grande Maison (The Big House) by Mohammed Dib [^5] https://en.wikipedia.org/wiki/La_Grande_Maison from French into Algerian Arabic (arq), and post-processed the translation to generate sentences for annotation by native speakers. Note that the translator is bilingual and a native speaker of Algerian Arabic.
  • News data: Although we prefer emotionally rich social media data from different platforms, when such data is scarce, we annotated news data and headlines in some African languages (e.g., yor, hau, and vmw).
  • Human-written and machine-generated data: We created a dataset from scratch for Hindi (hin) and Marathi (mar). Annotators were asked to come up with emotive sentences on a given topic (e.g., family). A small portion of the Hindi dataset was automatically translated into Marathi and manually corrected by native speakers to fix translation errors. Finally, we augmented both datasets with a few hundred quality-approved instances generated by ChatGPT. Note that these constitute less than 1% of the total number of data instances.
@@ 3.2 Data Annotation
Once the perceived emotion labels are assigned, the final intensity scores for Track B are determined by averaging the selected intensity values and rounding up to the nearest whole number. Intensity scores are assigned only for datasets in which most instances were annotated by at least five annotators to ensure robustness. Table 1 shows the total number of instances in each dataset, as well as the number of instances in the training, development, and test splits for all languages.
@@ 4.1 Tracks
Given a text and six emotion classes (i.e., joy, sadness, fear, anger, surprise, and disgust), participants were required to predict whether the intensity of each emotion was 0 (no emotion), 1 (low), 2 (medium), or 3 (high). Note that Track B does not include all languages, as intensity scores were only released for datasets with at least five annotators per instance to ensure more robust and reliable labels.
@@
Similar to Track A, participants were required to predict the presence or absence of each perceived emotion, but without using any training data in the target language. Instead, they were permitted to use labeled dataset(s) from at least one other language. For instance, one could use German data for training when testing on English. This track focuses on cross-lingual transfer and explores how data from various languages can support emotion detection in low-resource settings, as well as the ability of models to generalise across domains.
@@ 4.2 Task Organisation
$M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.232 $M_{\text{baseline}}$ 0.357 $M_{\text{baseline}}$ 0.243
@@
tir nta 0.591 ukr pai 0.726 vmw team unibuc 0.325 yor pai 0.461
late-gil-nlp 0.587 csiro-lt 0.664 pai 0.255 heimerdinger 0.392
$R_{\text{baseline}}$ 0.463 $R_{\text{baseline}}$ 0.535 $R_{\text{baseline}}$ 0.121 $R_{\text{baseline}}$ 0.092
$M_{\text{baseline}}$ 0.253 $M_{\text{baseline}}$ 0.157 $M_{\text{baseline}}$ 0.163 $M_{\text{baseline}}$ 0.165
Table 2: Average macro-F1 scores for our baselines ($M_{baseline}$ and $R_{baseline}$, referring to the Majority Vote and RoBERTa baselines, respectively) and the top two performing systems in Track A (shown in bold) for each language.
We used Codabench as the competition platform and released pilot datasets before the start of the shared task to help participants better understand the task (i.e., the datasets, the languages involved, and the labels). We provided participants with a starter kit on GitHub, resources for beginners, and organised a Q&A session along with a writing tutorial for junior researchers. Our participants were based in different parts of the world, as shown in Figure 2, with many coming from underrepresented regions. The task consisted of two phases: (1) the development phase and (2) the evaluation phase. During the development phase, the leaderboard was open, allowing a maximum of 999 submissions per participant. In the evaluation phase, the leaderboard was closed, and each participant was allowed up to three submissions, with the last submission being considered for the official ranking.
4.3 Evaluation Metrics and Baselines
Evaluation Metrics
Our Baselines
5 Participating Systems and Results
@@ 5.2.1 Best-Performing Systems
proposes one of the most effective models in the competition. They consistently rank as the top approach in Track A for 20 out of 28 languages. For their system, they combine several base models (ChatGPT-4o OpenAI (2024), DeepSeek-V3 DeepSeek-AI et al. (2025), Gemma-9b Team et al. (2024), Qwen-2.5-32b Yang et al. (2024), Mistral-Small-24B Jiang et al. (2024)) using multiple ensemble techniques (neural networks, XGBoost, LightGBM, linear regression, weighted voting). They fine-tune Gemma-9b and Qwen-2.5-32b using AdaLoRA. For prompting the LLMs, they used an iterative prompt-optimisation technique that generates prompt variations.
@@
ranks in the top 10 in 16 languages in Track A and 12th in English. They use sample contrastive learning, where performance is enhanced by comparing sample pairs, and generative contrastive learning, where the models learn to distinguish correct from incorrect predictions. Their samples are randomly selected from the task dataset (no external augmentation). They use LLaMa3-Instruct-8B AI@Meta (2024) for their fine-tuning.
@@ 5.2.2 Takeaways
Most of the teams that rank well on Track A experiment with essentially two methodologies: 1) fine-tuning BERT-based models such as DeBERTa Siino (2024), mBERT Dolev (2023), and XLM-R Conneau et al. (2020); and/or 2) instruction-fine-tuning using LoRa methodologies in combination with prompt design and data augmentation techniques on LLMs (ChatGPT-4o, DeepSeek-V3, Gemma-9b, Qwen-2.5-32b, Mistral-Small-24B). For instance, Team Telai (3rd in Chinese), Team Empaths (2nd in Hausa and Somali, and 1st in Swahili), Team NYCU-NLP (2nd in English), Team JNLP (1st in Hindi and among the best four across 10 languages), Team Unibouc (1st in Emakhuwa), Team Heiderdinger (2nd in Mozambican Portuguese, German, Spanish, Brazilian Portuguese, and Yorùbá; 1st in Russian), and Team Maomao (2nd in Afrikaans).
Few teams focus on only a subset of languages or explore language-related knowledge in their methodology. For instance, Team Lazarus NLP redefined and reformulated the multi-label classification into multiple binary tasks to expand training samples. They also explored how knowledge could potentially be transferred between Indonesian languages. All the top 10 teams performed significantly better than our baseline model, with an improvement that is more notable in a few low-resource languages, such as Oromo, where Team Tewodros obtained an average macro-F1 score of 0.616 compared to a baseline of 0.126. The same was observed for Yorùbá, where Team Pai scored 0.461 compared to a baseline score of 0.092.
@@ 5.3 Track B: Emotion Intensity Detection
$R_{\text{baseline}}$ 0.405 $R_{\text{baseline}}$ 0.562
@@
Table 3: Pearson correlation scores for our baselines (Majority: $M_{baseline}$ and RoBERTa: $R_{baseline}$) and the top two performing systems in Track B (shown in bold) for each language.
@@
$M_{\text{baseline}}$ 0.312 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.236 $M_{\text{baseline}}$ 0.254
@@
$M_{\text{baseline}}$ 0.204 $M_{\text{baseline}}$ 0.218 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.232
@@
$M_{\text{baseline}}$ 0.357 $M_{\text{baseline}}$ 0.243 $M_{\text{baseline}}$ 0.163 $M_{\text{baseline}}$ 0.652
@@
$M_{\text{baseline}}$ 0.162 $M_{\text{baseline}}$ 0.115 $M_{\text{baseline}}$ 0.165 $M_{\text{baseline}}$ 0.109
Table 4: Average macro-F1 scores for our baselines ($M_{baseline}$ and $R_{baseline}$, referring to the Majority Vote and RoBERTa baselines, respectively) and the top two performing systems in Track C (shown in bold) for each language.
5.3.1 Best-Performing Systems
Team Pai
Team CSECU-Learners
5.3.2 Takeaways
Teams Deepwave, Teleai, and JNLP also ranked highly across various languages using prompt engineering approaches similar to those in Track A. Additionally, Team NYCU-NLP ranked second in English by aggregating instruction-tuned small language models. All these teams outperformed our RoBERTa baseline, which achieved moderate Pearson correlation coefficient scores overall, but performed poorly in languages such as Algerian Arabic, Hausa, Ukrainian, and even Brazilian Portuguese -highlighting the difficulty of the task.
Overall, we observe that most teams adopted approaches similar to those used in Track A, with only minor adjustments to the prompts. Notably, even the best-performing teams achieved a Pearson correlation coefficient of no more than 0.65 on Algerian Arabic, likely due to the novelty and complexity of the dataset.
5.4 Track C: Cross-lingual Emotion Detection
Figure 3: Top LLMs used by participants across tracks (A, B, and C).
@@ 5.4.1 Best-Performing Systems
Team Deepwave fine-tuned Google Gemma-2 Team et al. (2024) using tailored data augmentation and Chain-of-Thought (CoT) prompting. They decomposed the task into two sub-tasks: (1) sentiment keyword identification and (2) sentiment polarity recognition.
@@
Team maomao experimented with different setups for fine-tuning LLMs. The base models used were Qwen2.5-7B-Instruct, GPT-0544o Mini2, and LLaMA-3.2-3B-Instruct (AI@Meta, 2024). They applied Direct Preference Optimisation to refine their model -a technique that selects high-quality instances from lower-quality ones within a dataset. After this step, they retrained the model using the refined dataset. They also explored random sampling and retrieval-augmented generation (RAG) methods for training, primarily on DeepSeek-V3, Qwen-Max5093, and Grok-V2.
@@ 5.4.2 Takeaways
Other top-performing systems include Team Ozemi, who fine-tuned a multilingual BERT model and applied machine translation to enhance performance across all languages. They used the Synthetic Minority Oversampling Technique (SMOTE) with TF-IDF to address class imbalance in Russian. They translated all the datasets into a common language using Google Translate before processing -except for Nigerian Pidgin and Emakhua, where they used a multilingual BERT model for translation. They also leveraged two Kaggle competition datasets for data augmentation.
Team heimerdinger, who ranked highest in Javanese, built their approach using various LLMs (LLaMA 3.1 8B, Qwen 2.5 7B, DeepSeek-7B, MistralV0.3-7B, and Gemma2-9B) for Track C. They employed in-context learning with multilingual examples from high-resource languages such as English and Spanish.
Overall, the participating teams outperformed our baselines. However, the average scores for this track are notably lower, particularly in low-resource languages, due to the additional challenges posed by limited data and resources. As shown in Table 4, there are significant performance gaps -even the top systems did not achieve an F-score higher than 0.50 in languages such as Javanese, Somali, Sundanese, Xhosa, Yorùbá, isiZulu, and Emakhuwa (where the top system achieved an F-score of only 0.21).